Circular No. 5884

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

ORIONID METEORS 1993 P. Brown, University of Western Ontario, writes: “Unusually high ac-

tivity has been reported from visual and photographic observers in Germany and the Netherlands of the Orionid meteor stream prior to its regular maximum. According to the International Meteor Organization (IMO), the regular maximum was predicted for Oct. 21.7 UT, while the outburst was recorded in the 2-hr interval centered around Oct. 18.1, with observers reporting uncorrected hourly rates of 25/hr. This corresponds to zenith hourly rates approaching 35/hr, some 3-4 times the regular activity of the Orionids at this solar longitude. The maximum Orionid ZHR over the past few years, according to the IMO, is ∼ 20.”

V1413 AQUILAE U. Munari, Asiago Astrophysical Observatory; and B. F. Yudin, Astro-

nomical Observatory, Moscow, write: “The eclipsing symbiotic nova V1413 Aql = AS 338 is entering a new outburst phase. Compared with similar observations made in June, échelle spectroscopy (range 450-700 nm) performed on Oct. 27.75 UT with the Asiago Observatory 1.82-m telescope (+ CCD) shows the disappearance of He I, He II, and [O III] emission lines. Only faint hydrogen Balmer lines remain visible. The star has been on a smooth decline since its 1982 outburst (cf. Munari 1992, A.Ap. 257, 163), and in 1990 it was observed at V = 13.0, B–V = +1.00, U–B = −0.20 (cf. Munari et al. 1992, A.Ap. Suppl. 93, 383). Photoelectric photometry performed on Oct. 27.85 by G. Sostero and A. Lepardo with the 0.45-m telescope of the Associazione Friulana di Astronomia e Meteorologia gives V = 11.2, B–V = +0.65.”

SUPERNOVA 1993J IN NGC 3031 J. A. Phillips and S. R. Kulkarni, California Institute of Technology,

report: “Using the five-element millimeter interferometer at the Owens Valley Radio Observatory, we have continued monitoring the flux density of SN 1993J at a frequency of 99.4 GHz: Sept. 12, 13 ± 2 mJy; Oct. 9, 8 ± 2 mJy. Our flux scale was established by observations of Neptune and Uranus. The 99-GHz flux density was nearly constant at 18 mJy from Apr. 11 until July 2 (IAUC 5763, 5775). The flux is now decaying and is indicative of optically thin expansion.”

1993 October 28 Daniel W. E. Green